OpenStudy (anonymous):

What is the 23rd term of the arithmetic sequence where a1 = 6 and a5 = 20?

OpenStudy (anonymous):

Can you define \( a_5\) in terms of \( a_1\) ?

OpenStudy (ash2326):

a1=6 first term for an Airthmetic Sequence nth term is given as an=a1+(n-1)*d where d is the common difference a1=6 a5=20 using a5 and a1 we can find d then use the nth term formula to find a23

OpenStudy (anonymous):

so do you agree that a5 = 6 +4d, therefore, it can be concluded that 6+4d =20 since a5 =20, therefore d = 7/2 To find a23 we now simply just sub in the no.s into the equation an=a1+(n-1)*d a23 =6+ (23-1)* 7/2 =83